OpenStudy (anonymous):

On the graph of the equation 3x + 2y = 18, what is the value of the y-intercept?

OpenStudy (anonymous):

-9?

OpenStudy (katrinakaif):

Make this a y = mx + b format equation 3x + 2y = 18 2y = -3x + 18 y = -3/2x + 9 Y-intercept is anything as "b" therefore, it is 9

OpenStudy (katrinakaif):

Close, but not quiet

OpenStudy (karatechopper):

9